OpenStudy (anonymous):

can someone help me for following question solve the following partial differential equation Using variable separable (u)_t=3u_xx 00 (u)_x (0,t)=u_x (2,t)=0 u(x,0)=3x

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (loser66):

\[\lambda^2 U_{xx} = U_t\\3U_{xx}=U_t\rightarrow \lambda^2 = 3\\U(x,0) = f(x) = 3x\\U(x,t) = \sum_{n=1}^\infty C_n e^{-n^2\pi^2\alpha^2t/L^2 }*sin\frac{n\pi x}{L}\] Is that what you need? if so, your L =6 and just find out \(C_n\) in U(x,t) which is \[C_n = \frac{2}{6}\int_0^6 3x sin \frac{n\pi x}{6}dx\]

OpenStudy (loser66):

@ybarrap hey, I forgot Fourier series formula for this problem. Can you check??

OpenStudy (ybarrap):

Is your problem $$ \cfrac{\partial u}{\partial t}=3\cfrac{\partial^2 u }{\partial x^2}\\ \cfrac{\partial u}{\partial x}\left ( 0,t\right )=\cfrac{\partial u}{\partial x}\left ( 2,t\right )=0\\ u(x,0)=3x $$ Where \( 0<x<2,~ t>0\)

OpenStudy (loser66):

he is not here , I just want to review what I learned. XD

OpenStudy (ybarrap):

The question is asking him to use "separation of variables" technique to solve.

OpenStudy (ybarrap):

I'm not sure how you would use Fourier series to solve, maybe Laplace, but not Fourier. Your equation for the coefficients looks right, but I don't see how to solve this problem using that technique.

OpenStudy (loser66):

ok, let mine aside, just show yours. hehee... I am crazy with this stuff.

OpenStudy (ybarrap):

ya!